OpenStudy (anonymous):

Which is the most reasonable estimate? forty-three percent of five hundred ten two hundred three hundred four hundred five hundred

OpenStudy (ranga):

round down 43 to 40 and 510 to 500. 40% of 500 = 500 * 40 / 100 = 5 * 40 = 200 is the estimate.

OpenStudy (anonymous):

thx